43. A coin is tossed four times. a) What is the probability of getting four heads? (1 mark) What is the probability of gettin and please can u show it that which part is a, b or c? tnx
Answer #1

a) The probability of getting 4 heads

n(s) = 2^4 = 16

A coin is toss means there is only one coin which is toss four times i.e.

= 1÷16

b) let the exactly two head be P(A)

Therefore, A = {HHTT , THHT, TTHH , HTHT, HTTH , THTH}

Therefore from classical definition we get P(A)

= 6÷16 or 3÷8

c) Probability of not getting two head I.e P( B)

P(B) = 1- P(A)

= 1 - ( 3÷ 8)

= 5 ÷ 8